OpenStudy (anonymous):

A Flower bed has the shape of a rectangle 5 yards long and 4 yards wide what is the area in square Feet ?

OpenStudy (anonymous):

@mathmate is the answer 180Ft^2

OpenStudy (theloshua):

correct

OpenStudy (mathmate):

Yes, it is correct. Either you do 5*4*9=180 sq. ft or (5*3)(4*3)=15*12=180 sq. ft.

OpenStudy (theloshua):

5 * 4 =20

OpenStudy (theloshua):

you could say that 5 yards = 15 feet ans that 4 yards = 12 feet

OpenStudy (theloshua):

15 8 12 = 180

OpenStudy (theloshua):

15* 20 = 180